SLOVE 6x 3/4 =7x 1/4 -2x 1/4

SLOVE 6x3/4=7x1/4-2x1/4

Grade:12

2 Answers

Ramvardhan
8 years ago
3/4
1/4
 
=> 6x3/4 = 5x1/4
=> 6x3/4 – 5x1/4 = 0
=> x1/4 ( 6x(3/4-1/4) - 5) = 0
=> x1/4 ( 6x1/2 - 5) = 0
=> x = 0, 25/36
 
 
charan
8 years ago
x = 0, 25/36